Please help me solve this:
Find the slope of the line 3y+2x=7

Answer by ewatrrr(24785)

 
Hi,
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
3y+2x=7 OR y = (-2/3)x + 7/2 m = (-2/3)